Please help me with this problem:
Two forces act on a point. One force is 38 newtons. The other force is 16 newtons. The angle between the two forces is 48°. Find the magnitude of the resultant force in newtons.
Enter answer accurate to four decimal places.
Thank you.

Answer by Alan3354(69443)

Two forces act on a point. One force is 38 newtons. The other force is 16 newtons. The angle between the two forces is 48°. Find the magnitude of the resultant force in newtons.
Enter answer accurate to four decimal places.
-------------------------
Make a parallelogram with the 2 vectors.
The resultant force is the long diagonal.
The angle opposite the diagonal is 180 - 48 = 132 degs.
--------
Use the Cosine Law:
r^2 = 38^2 + 16^2 - 2*38*16*cos(132)